shift summary for eve 6/29/99

Jim Graham (jagraham@hep.uchicago.edu)
Wed, 30 Jun 1999 00:36:38 -0500

Shift Summary for eve shift, 28 Jun 99

On Shift: JAG AE

Hours of beam delivered at all intensities: 7.5
Hours of beam delivered at nominal intensity: 7.5
Hours of E832 data (special runs not included): 6.0
Hours of special runs (laser scans, muon runs...): 0.3

Accounting of lost time: tape drive change, run change, analysis magnet
ramping

========================================================================
Runs ended this shift:
Run 13760 - 16.5M events - BEAM_832 sem=6.5E12
Run 13762 - 3.9M events - BEAM_832 sem=6.5E12
Run 13763 - 6.9M events - BEAM_832 sem=5.0E12
Run 13765 - - BEAM_MUON

Tapes brought to FCC this shift: None

========================================================================
Accesses during shift: None

========================================================================
Overview of shift:
Intensity studies at 6.5E12 continued ...
Ran at 5E12 for latter part of shift.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13760 16:00 inherit from day shift
13760 261 16:23 beam goes away, due to fuzzer problems in TeVatron
16:56 beam returns
17:24 end run to facilitate tape change. Requires taking
Ktev1 down.
Jason Ladue used opportunity to "massage" 2Y-0
banana input cables.
17:40 Dave Northacker puts 7 Ar_eth bottles "online"
18:00 Valeri Jejer updates fera peds in dbase.

Tape drive update; the replacement tape was installed
with the incorrect DIP switches set -> it is not
available to write to, and we will continue to write
data in the current configuration ( with the QKS
tape on Ktev1 in drive 5). This should get fixed
tomorrow day during some downtime.

13762 1 18:22 Start run
62 19:45 HCC (Et) bits for channels 248, 1703 are intermittently
inefficient. 248 is known, but I don't know about 1703.
19:51 End run to go down to 5E12

13763 1 19:55 start run, tuning down to 5E12
7 at 5E12
spooled out onsplit tape QKLA05. Checked and
deleted disk files for QKLA05, QKN005, QKNL05

modified BEAM_MUON config to use drive 5 on ktev1

~110 The Fera scaler in crate 53 raised a spurious
busy, hanging the trigger. Kfrend doesn't work,
since resetting this module would cost us the spill
number. We page experts (jason H, RSK) who tell us
we need to start a new run. Instead, we unplug the
Fera -command cable to this module, and reseat it,
causing it to drop REQ, and that seems to fix the
problem. We kfrend and continue, and Theo doesn't
complain about losing the spill number, so I think
all is well.

23:10 Tapes fill up, end run, go to muon mode.

13765 Muon run. Hogan instructs Miguel to raise DC HVs
for this run, which he does.
13766 Muon run with DCs back to nominal HV, (continues into
owl shift).

With help from Tsuyoshi, I edited
$DAMP_DIR/templates/.damprc.ktev (the Tcl job that
monitors the tape usage) to point at drive2 instead
of drive3, since drive2 fills up faster now. Should
take effect next fresh_start. The old one is in
$DAMP_DIR/templates/.damprc.ktev.Jun18_1998.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Tape drive 3 on Ktev1 was configured incorrectly. Needs to be
reconfigured in hardware tomorrow dayshift when we have a down time.

Regards, Jim ( for Albert)